The length of a rectangular carpet is 4 times its width. The area of the carpet is 144 square feet. What is the length of the carpet?.
Answer by jorel555(1290)

Area=Length x Width
144=x(4x)
4x^2=144
x^2=36
x=6
4x=24
The carpet is 24 x 6, so its' length is 24 feet!!!!!!!